# Bias and Fairness in Large Language Models: A Survey

### v2

@@ 3.5.2 Classifier Metrics
Classifier-based metrics rely on an auxiliary model to score generated text outputs for their toxicity, sentiment, or any other dimension of bias. Bias can be detected if text generated from similar prompts, but with different social groups, are classified differently. One prominent direction of research has been in toxicity detection. Perspective API [^3] https://perspectiveapi.com, developed by Google Jigsaw is a toxicity detection tool widely used in the literature (Bommasani et al., 2023; Chung et al., 2022; Chowdhery et al., 2022; Gehman et al., 2020). Given a text generation, Perspective API outputs a toxicity probability. For instance, to score generated text from any LLM, Gehman et al. (2020) leverage Expected Maximum Toxicity (EMT) computed over 25 generations and use mean and standard deviation. This metric characterizes the worst-case generations of an LLM. With Toxicity Probability (TP), they measure the empirical probability of generating at least one text with a toxicity score $\geq 0.5$ over the 25 generations, which captures how frequently the LLM generates text that is toxic. Other works have used the Toxic Fraction (TF), defined as the fraction of generations that are toxic (Bommasani et al., 2023). For a toxicity detector $c:\hat{Y}\rightarrow[0,1]$:
@@
Other methods have proposed more general approaches independent of Perspective API. Score Parity (Sicilia & Alikhani, 2023) measures how consistently a model generates language, as measured by a toxicity or sentiment classifier, given some protected attribute. For some scoring function $c:\hat{Y}\times A\rightarrow[0,1]$, lack of parity can be measured by:
@@
This notion can be generalized to demographic parity by using a different scoring function independent of the protected attribute. Counterfactual Sentiment Bias (Huang et al., 2020) similarly compares the sentiment of two sentences, generated with counterfactual prompts with a protected attribute replaced. This metric uses the Wasserstein-1 distance $\mathcal{W}_{1}(\cdot)$ between the sentiment distributions from some classifier $c:\hat{Y}\rightarrow[0,1]$:
@@
Classifier-based methods, however, need not be confined to sentiment or toxicity. For instance, Regard Score (Sheng et al., 2019) measures polarity towards and perceptions of social groups, similar to a sentiment and respect score. Using prefix templates to prompt the language model (e.g., "The woman worked as"), Sheng et al. (2019) generate a human-annotated dataset to train a regard classifier. Full Gen Bias (Smith et al., 2022) uses a style classifier to compute a style vector for each generated sentence $\hat{Y}_{w}$ corresponding to a term $w\in W$ in the prompt. Each element is the probability of a sentence belonging to one of $C$ style class, i.e., $\left[c(\hat{Y})[1],\cdots,c(\hat{Y})[C]\right]$. Bias is calculated as the variance across all generations:
@@ 3.5.3 Lexicon Metrics
Lexicon-based metrics perform a word-level analysis of the generated output, comparing each word to a pre-compiled list of harmful words, or assigning each word a pre-computed bias score. HONEST (Nozza et al., 2021) measures the number of hurtful completions. For identity-related template prompts and the top-$k$ completions $\hat{\mathbb{Y}}_{k}$, the metric calculates how many completions contain words in the HurtLex lexicon (Bassignana et al., 2018), given by:
@@ 3.5.4 Discussion and Limitations
We also discuss limitations of each class of generated text-based metrics. As Cabello et al. (2023) point out, word associations with protected attributes may be a poor proxy for downstream disparities, which may limit distribution-based metrics. Classifier-based metrics may be unreliable if the classifier itself has its own biases. For example, toxicity classifiers may disproportionately flag African-American English (Mozafari et al., 2020; Sap et al., 2019), and sentiment classifiers may incorrectly classify statements about stigmatized groups (e.g., people with disabilities, mental illness, or low socioeconomic status) as negative (Mei et al., 2023). Similarly, Pozzobon et al. (2023) highlight that automatic toxicity detection are not static and are constantly evolving. Thus, research relying solely on these scores for comparing models may result in inaccurate and misleading findings. These challenges may render classifier-based metrics themselves biased and unreliable. Finally, lexicon-based metrics may be overly coarse and overlook relational patterns between words, sentences, or phrases. Biased outputs can also be constructed from sequences of words that appear harmless individually, which lexicon-based metrics do not fully capture.
@@ 4 Taxonomy of Datasets for Bias Evaluation
In this section, we present datasets used in the literature for the evaluation of bias and unfairness in LLMs. We provide a taxonomy of datasets organized by their structure, such as masked sentences or sentence completions, which can guide metric selection. In Table 4, we summarize each dataset by the bias issue it addresses and the social groups it targets.
@@
Table 4: Taxonomy of Datasets for Bias Evaluation in LLMs. For each dataset, we show the number of instances in the dataset, the bias issue(s) they measure, and the group(s) they target. Black checks indicate explicitly stated issues or groups in the original work, while grey checks show additional use cases. For instances, while Winograd schema for bias evaluation assess gender-occupation stereotypes,(i) the stereotypes often illustrate a misrepresentation of gender roles, (ii) the model may have disparate performance for identifying male versus female pronouns, and (iii) defaulting to male pronouns, for example, reinforces exclusionary norms. Similarly, sentence completions intended to measure toxicity can trigger derogatory language.
@@
    Misrepresentation     Stereotyping     Disparate Performance     Derogatory Language     Exclusionary Norms     Toxicity     Age     Disability     Gender (Identity)     Nationality     Physical Appearance     Race     Religion     Sexual Orientation     Other
@@ 4.1 Counterfactual Inputs
Pairs or tuples of sentences can highlight differences in model predictions across social groups. Pairs are typically used to represent a counterfactual state, formed by perturbing a social group in a sentence while maintaining all other words and preserving the semantic meaning. A significant change in the model’s output – in the probabilities of predicted tokens, or in a generated continuation – can indicate bias.
@@ 4.1.1 Masked Tokens
One of the most prominent classes of these datasets is posed for coreference resolution tasks. The Winograd Schema Challenge was first introduced by Levesque et al. (2012) as an alternative to the Turing Test. Winograd schemas present two sentences, differing only in one or two words, and ask the reader (human or machine) to disambiguate the referent of a pronoun or possessive adjective, with a different answer for each of the two sentences. Winograd schemas have since been adapted for bias evaluation to measure words’ associations with social groups, most prominently with Winogender (Rudinger et al., 2018) and WinoBias (Zhao et al., 2018), with the form (with an example from Winogender):
@@
WinoBias measures stereotypical gendered associations with 3,160 sentences over 40 occupations. Some sentences require linking gendered pronouns to their stereotypically-associated occupation, while others require linking pronouns to an anti-stereotypical occupation; an unbiased model should perform both of these tasks with equal accuracy. Each sentence mentions an interaction between two occupations. Some sentences contain no syntactic signals (Type 1), while others are resolvable from syntactic information (Type 2). Winogender presents a similar schema for gender and occupation stereotypes, with 720 sentences over 60 occupations. While WinoBias only provides masculine and feminine pronoun genders, Winogender also includes a neutral option. Winogender also differs from WinoBias by only mentioning one occupation, which instead interacts with a participant, rather than another occupation. WinoBias+ (Vanmassenhove et al., 2021) augments WinoBias with gender-neutral alternatives, similar to Winogender’s neutral option, with 3,167 total instances.
@@
Though Winogender and WinoBias have been foundational to coreference resolution for bias evaluation, they are limited in their volume and diversity of syntax. Consequently, several works have sought to expand coreference resolution tests. GAP (Webster et al., 2018) introduces 8,908 ambiguous pronoun-name pairs for coreference resolution to measure gender bias. To represent more realistic use cases, this dataset is derived from Wikipedia. Not all examples follow Winograd schemas, but they all contain two names of the same gender and an ambiguous pronoun. The dataset contains an equal number of masculine and feminine instances. GAP-Subjective (Pant & Dadu, 2022) expands on GAP to include more subjective sentences expressing opinions and viewpoints. To construct the dataset, GAP sentences are mapped to a subjective variant (e.g., adding the word "unfortunately" or "controversial" to a sentence) using a style transfer model; thus, GAP-Subjective is the same size as GAP, with 8,908 instances. BUG (Levy et al., 2021) provides more syntactically diverse coreference templates, containing 108,419 sentences to measure stereotypical gender role assignments. The dataset is constructed by matching three corpora to 14 syntactic patterns that mention a human subject and referring pronoun, each annotated as stereotypical or anti-stereotypical.
@@
Other masked token datasets have been proposed for more general tasks, beyond coreference resolution. One of the most widely used is StereoSet (Nadeem et al., 2021), presented with the CAT metric (Section 3.4.2). StereoSet presents 16,995 crowdsourced instances measuring race, gender, religion, and profession stereotypes. For each type of bias, the dataset presents a context sentence with three options: one with a stereotype, one with a neutral or positive connotation (anti-stereotype), and one unrelated. StereoSet evaluates intrasentence bias within a sentence with fill-in-the-blank sentences, where the options describe a social group in the sentence context, such as:
@@
It measures intersentence bias between sentences in a discourse with three continuation options, where the first sentence mentions a social group. Providing similar sentences but without explicit options, Bias Evaluation Corpus with Professions (BEC-Pro) (Bartl et al., 2020) measures gender biases with respect to occupations, with 5,400 sentences constructed from templates containing a person word and one of 60 profession terms. For evaluation, person and profession words are masked.
@@ 4.1.2 Unmasked Sentences
Crowdsourced Stereotype Pairs (CrowS-Pairs) (Nangia et al., 2020) provides 1,508 sentence pairs to evaluate stereotypes of historically disadvantaged social groups. Each pair consists of one sentence demonstrating a stereotype, and the other sentence replacing the disadvantaged group with an advantaged one (anti-stereotype). The dataset covers nine types of biases: race, gender, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. For example:
@@
Several other sentence pair datasets follow similar forms. Equity Evaluation Corpus (Kiritchenko & Mohammad, 2018) contains 8,640 sentences to measure differences in sentiment towards gender and racial groups. The sentences are generated from templates instantiated with person and emotional state words, with tuples containing the same words except for the person term. RedditBias (Barikeri et al., 2021) introduces a conversational dataset generated from Reddit conversations to assess stereotypes between dominant and minoritized groups along the dimensions of gender, race, religion, and queerness. The dataset contains 11,873 sentences constructed by querying Reddit for comments that contain pre-specified sets of demographic and descriptor words, with human annotation to indicate the presence of negative stereotypes. To evaluate for bias, counterfactual sentence pairs are formed by replacing demographic terms with alternative groups. HolisticBias (Smith et al., 2022) contains 460,000 sentence prompts corresponding to 13 demographic axes with nearly 600 associated descriptor terms, generated with a participatory process with members of the social groups. Each sentence contains a demographic descriptor term in a conversational context, formed from sentence templates with inserted identity words. WinoQueer (Felkner et al., 2023) is a community-sourced dataset of 45,540 sentence pairs to measure anti-LGBTQ+ stereotypes, curated by surveying members of the LGBTQ+ community. Each pair contains a sentence mentioning a LGBTQ+ identity descriptor, and a counterfactual version with a non-LGBTQ+ identity. Bias-STS-B (Webster et al., 2020) adapts the original Semantic Textual Similarity-Benchmark (STS-B) (Cer et al., 2017) to generate pairs of sentences differing only in gendered terms, but otherwise maintaining the same meaning for sentences in a pair. PANDA (Qian et al., 2022) introduces a dataset of 98,583 text perturbations for gender, race/ethnicity, and age groups, with pairs of sentences with a social group changed but the semantic meaning preserved. PANDA includes annotations for the perturbed demographic words. Though originally proposed as a dataset for fine-tuning, the dataset can also be used to assess robustness to demographic perturbation, where a fair model produces two invariant outputs given an input sentence and its perturbation.
Bias NLI (Dev et al., 2020) alternatively probes for bias through inference tasks. The dataset presents sentence pairs that should not entail nor contradict each other, but may be predicted to imply one another if representational biases are present in a model. The dataset instantiates the template "the [subject] [verb] a/an [object]", instantiated with gender-occupation, nationality-polarity (e.g., "evil person", "good person"), and religion-polarity subject pairs. For example:

### v3

@@ 3.5.2 Classifier Metrics
Classifier-based metrics rely on an auxiliary model to score generated text outputs for their toxicity, sentiment, or any other dimension of bias. Bias can be detected if text generated from similar prompts, but with different social groups, are classified differently. One prominent direction of research has been in toxicity detection. Perspective API [^3] https://perspectiveapi.com, developed by Google Jigsaw is a toxicity detection tool widely used in the literature (Liang et al., 2022; Chung et al., 2022; Chowdhery et al., 2022; Gehman et al., 2020). Given a text generation, Perspective API outputs a toxicity probability. For instance, to score generated text from any LLM, Gehman et al. (2020) leverage Expected Maximum Toxicity (EMT) computed over 25 generations and use mean and standard deviation. This metric characterizes the worst-case generations of an LLM. With Toxicity Probability (TP), they measure the empirical probability of generating at least one text with a toxicity score $\geq 0.5$ over the 25 generations, which captures how frequently the LLM generates text that is toxic. Other works have used the Toxic Fraction (TF), defined as the fraction of generations that are toxic (Liang et al., 2022). For a toxicity detector $c:\hat{Y}\rightarrow[0,1]$:
@@
Other methods have proposed more general approaches independent of Perspective API. Score Parity (Sicilia and Alikhani, 2023) measures how consistently a model generates language, as measured by a toxicity or sentiment classifier, given some protected attribute. For some scoring function $c:\hat{Y}\times A\rightarrow[0,1]$, lack of parity can be measured by:
@@
Counterfactual Sentiment Bias (Huang et al., 2020) similarly compares the sentiment of two sentences, generated with counterfactual prompts with a protected attribute replaced. This metric uses the Wasserstein-1 distance $\mathcal{W}_{1}(\cdot)$ between the sentiment distributions from some classifier $c:\hat{Y}\rightarrow[0,1]$:
@@
Classifier-based methods, however, need not be confined to sentiment or toxicity. For instance, Regard Score (Sheng et al., 2019) measures polarity towards and perceptions of social groups, similar to a sentiment and respect score. Using prefix templates to prompt the language model (e.g., "The woman worked as"), Sheng et al. (2019) generate a human-annotated dataset to train a regard classifier. Full Gen Bias (Smith et al., 2022) uses a style classifier to compute a style vector for each generated sentence $\hat{Y}_{w}$ corresponding to a term $w\in W$ in the prompt. Each element is the probability of a sentence belonging to one of $C$ style class, namely, $\left[c(\hat{Y})[1],\cdots,c(\hat{Y})[C]\right]$. Bias is calculated as the variance across all generations:
@@ 3.5.3 Lexicon Metrics
Lexicon-based metrics perform a word-level analysis of the generated output, comparing each word to a pre-compiled list of harmful words, or assigning each word a pre-computed bias score. HONEST (Nozza, Bianchi, and Hovy, 2021) measures the number of hurtful completions. For identity-related template prompts and the top-$k$ completions $\hat{\mathbb{Y}}_{k}$, the metric calculates how many completions contain words in the HurtLex lexicon (Bassignana et al., 2018), given by:
@@ 3.5.4 Discussion and Limitations
We also discuss the limitations of each class of generated text-based metrics. As Cabello, Jørgensen, and Søgaard (2023) point out, word associations with protected attributes may be a poor proxy for downstream disparities, which may limit distribution-based metrics that rely on vectors of co-occurrence counts. For example, co-occurrence does not account for use-mention distinctions, where harmful words may be mentioned in the same context of a social group (e.g., as counterspeech) without using them to target that group (Gligoric et al., 2024). Classifier-based metrics may be unreliable if the classifier itself has its own biases. For example, toxicity classifiers may disproportionately flag African-American English (Mozafari, Farahbakhsh, and Crespi, 2020; Sap et al., 2019), and sentiment classifiers may incorrectly classify statements about stigmatized groups (e.g., people with disabilities, mental illness, or low socioeconomic status) as negative (Mei, Fereidooni, and Caliskan, 2023). Similarly, Pozzobon et al. (2023) highlight that automatic toxicity detection are not static and are constantly evolving. Thus, research relying solely on these scores for comparing models may result in inaccurate and misleading findings. These challenges may render classifier-based metrics themselves biased and unreliable. Finally, lexicon-based metrics may be overly coarse and overlook relational patterns between words, sentences, or phrases. Biased outputs can also be constructed from sequences of words that appear harmless individually, which lexicon-based metrics do not fully capture.
3.6 Recommendations
We synthesize findings and guidance from the literature to make the following recommendations. For more detailed discussion and limitations, see Sections 3.3.3, 3.4.3, and 3.5.4.
  1. 1.
    Exercise caution with embedding-based and probability-based metrics. Bias in the embedding space can have a weak and unreliable relationship with bias in the downstream application. Probability-based metrics also show weak correlations with downstream biases. Therefore, embedding- and probability-based metrics should be avoided as the sole metric to measure bias and should instead be accompanied by a specific evaluation of the downstream task directly.
  2. 2.
    Report model specifications. The choice of model hyperparameters can lead to contradictory conclusions about the degree of bias in a model. Bias evaluation should be accompanied by the model specification and the specific templates or prompts used in calculating the bias metric.
  3. 3.
    Construct metrics to reflect real-world power dynamics. Nearly all metrics presented here employ some notion of invariance, via Definitions 9, 10, 11, or 12 in Section 2.3. Differences in linguistic associations can encode important, non-stereotypical knowledge about social groups, so usage of these metrics should explicitly state the targeted harm. Metrics that rely on auxiliary datasets or classifiers, particularly pseudo-log-likelihood and classifier metrics, should ensure that the auxiliary resource measures the targeted bias with construct and ecological validity.
Given the limitations of the existing metrics, it may be necessary to develop new evaluation strategies that are explicitly and theoretically grounded in the sociolinguistic mechanism of bias the metric seeks to measure. In constructing new metrics, we reiterate Cao et al. (2022b)’s desiderata for measuring stereotypes, which can be extended to other forms of bias: (1) natural generalization to previously unconsidered groups; (2) grounding in social science theory; (3) exhaustive coverage of possible stereotypes (or other biases); (4) natural text inputs to the model; and (5) specific, as opposed to abstract, instances of stereotypes (or other biases).
@@ 4 Taxonomy of Datasets for Bias Evaluation
In this section, we present datasets used in the literature for the evaluation of bias and unfairness in LLMs. We provide a taxonomy of datasets organized by their structure, which can guide metric selection. In Table 4, we summarize each dataset by the bias issue it addresses and the social groups it targets.
@@
Table 4: Taxonomy of Datasets for Bias Evaluation in LLMs. For each dataset, we show the number of instances in the dataset, the bias issue(s) they measure, and the group(s) they target. Black checks indicate explicitly stated issues or groups in the original work, while grey checks show additional use cases. For instance, while Winograd schema for bias evaluation assess gender-occupation stereotypes, (i) the stereotypes often illustrate a misrepresentation of gender roles, (ii) the model may have disparate performance for identifying male versus female pronouns, and (iii) defaulting to male pronouns, for example, reinforces exclusionary norms. Similarly, sentence completions intended to measure toxicity can trigger derogatory language.
@@
    Misrepresentation     Stereotyping     Disparate Performance     Derogatory Language     Exclusionary Norms     Toxicity     Age     Disability     Gender (Identity)     Nationality     Physical Appearance     Race     Religion     Sexual Orientation     $\textrm{}^{\dagger}$
@@
$\textrm{}^{\dagger}$xamples of other social axes include socioeconomic status, political ideology, profession, and culture.
@@ 4.1 Counterfactual Inputs
Pairs or tuples of sentences can highlight differences in model predictions across social groups. Pairs are typically used to represent a counterfactual state, formed by perturbing a social group in a sentence while maintaining all other words and preserving the semantic meaning. A significant change in the model’s output — in the probabilities of predicted tokens, or in a generated continuation — can indicate bias.
@@ 4.1.1 Masked Tokens
One of the most prominent classes of these datasets is posed for coreference resolution tasks. The Winograd Schema Challenge was first introduced by Levesque, Davis, and Morgenstern (2012) as an alternative to the Turing Test. Winograd schemas present two sentences, differing only in one or two words, and ask the reader (human or machine) to disambiguate the referent of a pronoun or possessive adjective, with a different answer for each of the two sentences. Winograd schemas have since been adapted for bias evaluation to measure words’ associations with social groups, most prominently with Winogender (Rudinger et al., 2018) and WinoBias (Zhao et al., 2018), with the form (with an example from Winogender):
@@
where [MASK] may be replaced by she, he, or they. WinoBias measures stereotypical gendered associations with 3,160 sentences over 40 occupations. Some sentences require linking gendered pronouns to their stereotypically-associated occupation, while others require linking pronouns to an anti-stereotypical occupation; an unbiased model should perform both of these tasks with equal accuracy. Each sentence mentions an interaction between two occupations. Some sentences contain no syntactic signals (Type 1), while others are resolvable from syntactic information (Type 2). Winogender presents a similar schema for gender and occupation stereotypes, with 720 sentences over 60 occupations. While WinoBias only provides masculine and feminine pronoun genders, Winogender also includes a neutral option. Winogender also differs from WinoBias by only mentioning one occupation, which instead interacts with a participant, rather than another occupation. WinoBias+ (Vanmassenhove, Emmery, and Shterionov, 2021) augments WinoBias with gender-neutral alternatives, similar to Winogender’s neutral option, with 3,167 total instances.
@@
Though Winogender and WinoBias have been foundational to coreference resolution for bias evaluation, they are limited in their volume and diversity of syntax. Consequently, several works have sought to expand coreference resolution tests. GAP (Webster et al., 2018) introduces 8,908 ambiguous pronoun-name pairs for coreference resolution to measure gender bias. To represent more realistic use cases, this dataset is derived from Wikipedia. Not all examples follow Winograd schemas, but they all contain two names of the same gender and an ambiguous pronoun. The dataset contains an equal number of masculine and feminine instances. GAP-Subjective (Pant and Dadu, 2022) expands on GAP to include more subjective sentences expressing opinions and viewpoints. To construct the dataset, GAP sentences are mapped to a subjective variant (e.g., adding the word "unfortunately" or "controversial" to a sentence) using a style transfer model; thus, GAP-Subjective is the same size as GAP, with 8,908 instances. BUG (Levy, Lazar, and Stanovsky, 2021) provides more syntactically diverse coreference templates, containing 108,419 sentences to measure stereotypical gender role assignments. The dataset is constructed by matching three corpora to 14 syntactic patterns that mention a human subject and referring pronoun, each annotated as stereotypical or anti-stereotypical.
@@
Other masked token datasets have been proposed for more general tasks, beyond coreference resolution. One of the most widely used is StereoSet (Nadeem, Bethke, and Reddy, 2021), presented with the CAT metric (Section 3.4.2). StereoSet presents 16,995 crowdsourced instances measuring race, gender, religion, and profession stereotypes. For each type of bias, the dataset presents a context sentence with three options: one with a stereotype, one with a neutral or positive connotation ("anti-stereotype"), and one unrelated. StereoSet evaluates intrasentence bias within a sentence with fill-in-the-blank sentences, where the options describe a social group in the sentence context, such as:
@@
It measures intersentence bias between sentences in a discourse with three continuation options, where the first sentence mentions a social group. Providing similar sentences but without explicit options, Bias Evaluation Corpus with Professions (BEC-Pro) (Bartl, Nissim, and Gatt, 2020) measures gender biases with respect to occupations, with 5,400 sentences constructed from templates containing a person word and one of 60 profession terms. For evaluation, person and profession words are masked.
@@ 4.1.2 Unmasked Sentences
Crowdsourced Stereotype Pairs (CrowS-Pairs) (Nangia et al., 2020) provides 1,508 sentence pairs to evaluate stereotypes of historically disadvantaged social groups. Each pair consists of one sentence demonstrating a stereotype, and the other sentence replacing the disadvantaged group with an advantaged one ("anti-stereotype"). The dataset covers nine types of biases: race, gender, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. For example:
@@
Several other sentence pair datasets follow similar forms. Equity Evaluation Corpus (Kiritchenko and Mohammad, 2018) contains 8,640 sentences to measure differences in sentiment towards gender and racial groups. The sentences are generated from templates instantiated with person and emotional state words, with tuples containing the same words except for the person term. RedditBias (Barikeri et al., 2021) introduces a conversational dataset generated from Reddit conversations to assess stereotypes between dominant and minoritized groups along the dimensions of gender, race, religion, and queerness. The dataset contains 11,873 sentences constructed by querying Reddit for comments that contain pre-specified sets of demographic and descriptor words, with human annotation to indicate the presence of negative stereotypes. To evaluate for bias, counterfactual sentence pairs are formed by replacing demographic terms with alternative groups. HolisticBias (Smith et al., 2022) contains 460,000 sentence prompts corresponding to 13 demographic axes with nearly 600 associated descriptor terms, generated with a participatory process with members of the social groups. Each sentence contains a demographic descriptor term in a conversational context, formed from sentence templates with inserted identity words. WinoQueer (Felkner et al., 2023) is a community-sourced dataset of 45,540 sentence pairs to measure anti-LGBTQ+ stereotypes, curated by surveying members of the LGBTQ+ community. Each pair contains a sentence mentioning a LGBTQ+ identity descriptor, and a counterfactual version with a non-LGBTQ+ identity. Bias-STS-B (Webster et al., 2020) adapts the original Semantic Textual Similarity-Benchmark (STS-B) (Cer et al., 2017) to generate pairs of sentences differing only in gendered terms, but otherwise maintaining the same meaning for sentences in a pair. PANDA (Qian et al., 2022) introduces a dataset of 98,583 text perturbations for gender, race/ethnicity, and age groups, with pairs of sentences with a social group changed but the semantic meaning preserved. PANDA includes annotations for the perturbed demographic words. Though originally proposed for fine-tuning, the dataset can also be used to assess robustness to demographic perturbation, where a fair model produces two invariant outputs given an input sentence and its perturbation.
Bias NLI (Dev et al., 2020) alternatively probes for bias through inference tasks. The dataset presents sentence pairs that should not entail nor contradict each other, but may be predicted to imply one another if representational biases are present in a model. The dataset instantiates the template "the [subject] [verb] a/an [object]", instantiated with gender-occupation, nationality-polarity (e.g., "evil person," "good person"), and religion-polarity subject pairs. For example: